3p(3p-2)=0
So, P=?

Answer by jim_thompson5910(35256)

3p(3p-2)=0


3p = 0 or 3p-2 = 0


3p = 0 or 3p = 2


p = 0/3 or p = 2/3


p = 0 or p = 2/3